OpenStudy (anonymous):

What is the solution set of |2x - 3| - 6 = -1? Choose one answer. a. 5 or -5 b. 4 or 1 c. 4 or -1 d. No solution

OpenStudy (anonymous):

c

OpenStudy (anonymous):

x = -1, 4

OpenStudy (anonymous):

|2x - 3| = 5 so 2x - 3 = +- 5 so 2x = 3 +- 5 so x = (3 +- 5)/2 so x = 4 or -1

OpenStudy (anonymous):

oh my so many different answer!

OpenStudy (anonymous):

they're all the same answer.

OpenStudy (anonymous):

haha i know i got confuzzled for a second :p